OpenStudy (anonymous):

What exponential function represents the data in the table? x f(x) 2 25 3 125 4 625

OpenStudy (anonymous):

not much of a table.... but yeah... here:

OpenStudy (anonymous):

f(x)=5^x f(2)=5^2=25 f(3)=5^3=125 f(4)=5^4=625

OpenStudy (anonymous):

so i graph them and see???

OpenStudy (anonymous):

nope u have x and f(x) put the values of x in each choice and see it gives the values of f(x)

OpenStudy (anonymous):

actually ooooooohhhhhhh ok... i just saw....

OpenStudy (anonymous):

sorry about that.... lol i'm basically blind

OpenStudy (anonymous):

literally blind ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

i'm nearsighted, and i can see stuff, but then again i don't see stuff

OpenStudy (anonymous):

wear glasses, I wear them 24/7

OpenStudy (anonymous):

like my mom will be like "get me that water bottle" and i'm like "what water bottle? Where" and she's like "RIGHT IN FRONT OF YOU" lol

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

and thanks so so so so much :D

OpenStudy (anonymous):

u r welcome :D

OpenStudy (anonymous):

@31356

OpenStudy (31356):

Correct! :D